OpenStudy (anonymous):

\sum _{n=1}^∞\left(7-b\right)=2 cal 3 anyone?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(7-b)^n=2\]

OpenStudy (anonymous):

I gotta find b

OpenStudy (anonymous):

Idk how to solve but I'm commenting to find out how later on

OpenStudy (michele_laino):

such series converges if and only if the subsequent condition holds: \[\left| {7 - b} \right| < 1\; \Rightarrow 6 < b < 8\] and the value of the corresponding sum is: \[\frac{1}{{1 - \left( {7 - b} \right)}}\] so, we have to solve this equation, for \(b\): \[\frac{1}{{1 - \left( {7 - b} \right)}} = 2\]

OpenStudy (kainui):

The key to this is knowing what this sum is equal to: \[S=\sum_{n=1}^\infty x^n\] and then plugging in your value of (7-b)=x into that. It'll hopefully become clear soon what I mean. So let's write that sum above, S without hiding the terms: \[S= x+x^2+x^3+\cdots\] notice, if we multiply the whole thing by x we get: \[xS=x^2+x^3+x^4+\cdots\] Now if we add x to it, what do we get? We get the same sum! \[xS+x= x+x^2+x^3+\cdots\]\[xS+x=S\] So with a few steps of algebra of solving for S we get an alternate way of writing S, \[S = \frac{x}{1-x}=\sum_{n=1}^\infty x^n\] Now you can plug in stuff and solve hopefully.

OpenStudy (michele_laino):

oops.. I have made an error I'm very sorry @Ratedover Here is the right equation: \[\Large \frac{{7 - b}}{{1 - \left( {7 - b} \right)}} = 2\] thanks!! for your reply :) @Kainui

OpenStudy (kainui):

Heh yeah that pesky index starting at 0 instead of 1, I figured I'd just go ahead and explain it cause of that :P

OpenStudy (michele_laino):

yes! well done !! :) @Kainui